OpenStudy (anonymous):

y = log x If y = 10, then what is x?

OpenStudy (anonymous):

y = log x 10 = log x 10 = 10^x x = 1

zepdrix (zepdrix):

\[\Large\bf\sf y=\log_{10}x\qquad\to\qquad 10^y=x\] When y is 10, \[\Large\bf\sf 10^{10}=x\]x ends up being a really really big number.

OpenStudy (anonymous):

hows that right? ^ its not log 10 x or 10 to the y power??

zepdrix (zepdrix):